A Complete Bibliography of Publications in Reviews of Modern Physics (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

25 April 2024 Version 1.07

Title word cross-reference

4 [NO20]. _{2} [SSN^{+}23]. _{3}^{+} [MTGS20]. b [BSRW22]. ββ [ABD^{+}23]. D [EH20].

-hadron [BSRW22].

2D [WKVT24].

90 [GHM^{+}22]. 92 [TH22]. 94 [HPF^{+}24].

accelerator [KJPW23]. accelerators [HKP^{+}22]. Achievement [Hal20]. actinide [CLvdL23]. additive [ZSC^{+}22]. Advances [EKTvR20, ZT23]. advantage [HE23]. aerosols [PPK^{+}23]. algorithms [BCLK^{+}22]. Angle [SHS21]. Angle-resolved [SHS21]. annihilation [ADK^{+}21]. Anomalous [ADKB23, CLM23]. application [CLvdL23]. applications [BMMS23]. Applied [TT22]. approaches [BHMV21, LLM^{+}20]. APS [Hal20]. ARPES [BZD24]. assembly [HG21]. associated [Var20]. asteroseismology [Aer21]. astronomy [MTGS20]. astrophysical [BBD^{+}21]. Atmospheric [SCB^{+}23]. Atom [KJPW23]. atoms [ADKB23, PLH^{+}24, Wu21]. Atomtronic [AAB^{+}22]. Auger [GOP^{+}21]. automation [ZT23]. axion [Sik21].

bacteria [MGF24]. Baryogenesis [BB21]. Be [Kam23]. bed [ZSC^{+}22]. behavior [BB20]. Bell [WA20]. between [RTK22]. beyond [CGPS22]. biological [MHHT20, TMLS22]. biomedical [BMMS23]. blowing [LSZ^{+}20]. body [AAB^{+}22, BMV^{+}21, Sch22]. bosons [AST21]. boundary [LPS22]. boundary-driven [LPS22]. brain [PD22]. Bunch [HKP^{+}22]. bundle [CGFA22]. bursts [Zha23].

capture [BED^{+}20, CSL^{+}21]. Case [Jag21]. Cavity [Rei22]. Cavity-enhanced [Rei22]. CeCu [SSN^{+}23]. cells [TMLS22]. challenges [EKTvR20]. change [GL20]. charge [BG23, GV22]. Charged [GBMB22]. chemistry [MEAG^{+}20]. Circuit [BGGW21]. circuits [AAB^{+}22]. classical [LP21, TH20, TH22, ZLM^{+}23]. climate [CS21, GL20, SO22, WTC^{+}21]. CODATA [TMNT21]. Collective [SPI^{+}23, WHK^{+}22, MGF24]. colliders [SZ21]. collisions [AST21, KNS20]. Colloquia [BK21]. Colloquium [ADKB23, AAB^{+}22, BG23, BLFV^{+}21, BMMS23, BBG^{+}20, BDS^{+}22, BEG^{+}23, CLM23, CGFA22, GR24, GHK^{+}23, HM22, KPC^{+}20, LLM^{+}20, MGF24, MA20, NO20, PK21, Pic23, PD22, Rei22, SS20, SBPM24, SSN^{+}23, Sun23, TH20, Var20, WKVT24, WA20, ZLM^{+}23, ZT23, dlTKC^{+}21, TH22]. Communications [Tho20]. Complex [ABBW22]. components [VTR20]. Comprehensive [PLH^{+}24]. Computational [HE23, MEAG^{+}20]. computing [MSUY22]. Concepts [CPGSV21]. condensed [Hal20, PK21, RTK22]. confinement [HPB^{+}23]. conjecture [HHRR23]. connections [BHMV21]. constants [TMNT21]. constraints [WTC^{+}21]. contextuality [BCG^{+}22]. control [FRG^{+}24, ZT23, dlTKC^{+}21]. Controlling [YZX^{+}24]. convection [SS20]. cooled [ADKB23]. correlated [VPC^{+}23]. correlations [SPI^{+}23]. counting [KJPW23]. Coupling [TMLS22, TH20, TH22]. creation [ABD^{+}23]. criticality [KPC^{+}20]. cryptography [PR22]. crystallizations [NO20]. crystals [ZLM^{+}23]. Culinary [MLPM23]. current [FRG^{+}24]. current-induced [FRG^{+}24]. currents [MLPM23].

Dark [ADK^{+}21]. decay [ABD^{+}23]. decays [BSRW22]. decentralized [WHK^{+}22]. defects [BB20]. deformation [TMLS22]. density [BG23, SBB^{+}23]. dependent [LLM^{+}20]. design [Pic23]. detectors [BBG^{+}20]. devices [XMZ^{+}20, ZT23]. diagnostics [HPF^{+}22, HPF^{+}24]. diamond [BSB^{+}20]. diffraction [FML^{+}22]. diffusion [YZX^{+}24]. dimensional [BHMK^{+}21, LQD21]. dimensions [Jag21]. discovery [ABD^{+}23]. discrete [ZLM^{+}23]. diseases [PPK^{+}23]. disorder [RRM22]. disordered [VPC^{+}23]. dissipation [HM22]. dissipative [ADKB23]. distribution [XMZ^{+}20]. dot [ZT23]. double [BED^{+}20]. double-electron [BED^{+}20]. driven [BDK^{+}23, LPS22]. droplets [PPK^{+}23]. dynamic [FML^{+}22]. Dynamics [GOP^{+}21, RRM22, Sch22, SS20, WHK^{+}22]. Dzyaloshinskii [KCS^{+}23].

Earth [VTR20]. Editorial [BK21, Kam23, Tho20, TT22]. effect [CLM23]. effective [HKvK20, IWW24, JLL^{+}21]. effects [MGF24]. Einstein [EH20]. elastic [BDK^{+}23]. Elasticity [BB20]. electric [FRG^{+}24]. Electrical [FRG^{+}24]. electrified [LSZ^{+}20]. electrodynamics [BGGW21, SPI^{+}23, SN22]. electromagnetic [GBMB22, TMLS22]. electron [BED^{+}20, EKTvR20, FML^{+}22, HKP^{+}22, KPC^{+}20]. electronic [PZLCEF^{+}22]. elements [CSL^{+}21]. Emergent [WTC^{+}21]. energy [KNS20, MA20, SBB^{+}23, YZX^{+}24]. engineered [SBPM24]. Engineering [PZLCEF^{+}22, FKK^{+}23]. enhanced [Rei22]. entangled [CPGSV21]. entropy [AHM^{+}21, LP21]. equations [EH20]. equilibria [Par23]. Equilibrium [HG21, NO20]. Erratum [GHM^{+}22, HPF^{+}24, TH22]. error [CBB^{+}23]. estimation [CM22]. Evolution [OGS^{+}20]. Exceptional [BBK21, Hal20]. exotic [OGS^{+}20]. Experimental [LQD21]. experiments [LCPP22].

factors [BEG^{+}23]. far [NO20]. fast [Zha23]. feature [FKK^{+}23]. feedbacks [CS21]. Fermi [CGPS22]. fermion [SSN^{+}23]. fiber [CGFA22]. Fibonacci [Jag21]. field [BMV^{+}21, DPV^{+}20, FRG^{+}24, HKvK20, IWW24, MSUY22, MA20]. fields [GBMB22, TMLS22]. films [KCS^{+}23]. Finite [BHMK^{+}21]. Finite-temperature [BHMK^{+}21]. flagellar [MGF24]. flavor [BSRW22]. flight [Sun23]. fluid [MLPM23]. food [MLPM23]. form [BEG^{+}23]. forty [Gen22]. forty-year [Gen22]. Fracton [GR24]. freedom [SN22]. frustration [BDK^{+}23]. fully [SSN^{+}23]. functional [Sch22]. fundamental [TMNT21]. fusion [HPB^{+}23, ZSC^{+}22]. future [CXZL22, SZ21].

gapped [SSN^{+}23]. gauge [SN22]. Geometric [CGFA22]. geometrical [BDK^{+}23]. Grand [VTR20, BB21]. Gravitational [BEG^{+}23, CM22]. gravity [BBD^{+}21, HHRR23]. grids [WHK^{+}22]. group [BLFV^{+}21]. group-IV [BLFV^{+}21]. growth [SCB^{+}23].

H [MTGS20]. hadron [BSRW22]. Hadronic [GHM^{+}18, KNS20, GHM^{+}22]. Hall [CLM23]. Hartree [LLM^{+}20]. Hawking [AHM^{+}21]. heat [BMV^{+}21, PK21]. heaviest [CSL^{+}21]. heavy [SSN^{+}23, KPC^{+}20]. Heavy-electron [KPC^{+}20]. heavy-fermion [SSN^{+}23]. Helfrich [BDK^{+}23]. helium [BB20]. Hermitian [BBK21]. hidden [Jag21]. high [KNS20, SBB^{+}23, Var20]. high-energy [KNS20]. high-energy-density [SBB^{+}23]. holography [BG23]. Home [TT22]. Hurault [BDK^{+}23]. Hydrodynamics [BG23]. identity [HLE^{+}22]. imaging [SBB^{+}23]. Implications [SN22]. including [Var20]. Inclusive [Tho20]. Incompatible [GHK^{+}23]. indistinguishability [HLE^{+}22]. indistinguishable [LLM^{+}20]. induced [FRG^{+}24]. inertial [HPB^{+}23]. infectious [PPK^{+}23]. information [GHK^{+}23]. initio [BHMV21]. insect [Sun23]. Insights [CGFA22]. instabilities [BDK^{+}23]. interacting [DDM^{+}23, RRM22]. interaction [KCS^{+}23]. interactions [SKS^{+}23, Wu21]. interdisciplinary [BHMV21]. interfaces [WKVT24]. Interfacial [CXZL22, KCS^{+}23]. interior [Aer21]. intermediate [BCLK^{+}22]. intermediate-scale [BCLK^{+}22]. internet [AEE^{+}23]. Interplay [RTK22]. Invisible [Sik21]. ions [MCD^{+}21]. Irreversible [LP21]. IV [BLFV^{+}21].

jets [AST21, LSZ^{+}20]. journey [Gen22].

key [XMZ^{+}20]. Kinematic [FKK^{+}23]. Kitaev [CGPS22]. Kochen [BCG^{+}22].

laboratory [BSRW22, SBB^{+}23]. Lamb [PLH^{+}24]. landscapes [PZLCEF^{+}22]. lapse [CS21]. Large [EH20, JLL^{+}21]. Large-momentum [JLL^{+}21]. Laser [ZSC^{+}22, ADKB23, HPF^{+}22, HPF^{+}24]. laser-cooled [ADKB23]. laser-produced [HPF^{+}22, HPF^{+}24]. lattice [BHMK^{+}21, MSUY22]. lattices [ADKB23]. Law [SO22]. learning [BDS^{+}22]. Lecture [Par23, Gen22]. lepton [BSRW22]. Light [VPC^{+}23, Wit20, CGFA22, PLH^{+}24]. limit [EH20]. limiting [HG21]. limits [MA20]. Linear [Var20, HKP^{+}22]. liquids [CGPS22]. locally [WA20]. Long [DDM^{+}23]. Long-range [DDM^{+}23]. lubric [WKVT24].

Machine [BDS^{+}22]. machines [MHHT20]. magnetic [KCS^{+}23, MA20, SKS^{+}23]. magnetism [FRG^{+}24]. magnetometry [BSB^{+}20]. Magnetotactic [MGF24]. manufacturing [ZSC^{+}22]. many [AAB^{+}22, BMV^{+}21, Sch22]. many-body [AAB^{+}22, BMV^{+}21, Sch22]. mass [KJPW23, YZX^{+}24]. material [WKVT24]. materials [BZD24, CLvdL23, Pic23, SHS21, dlTKC^{+}21]. Matrix [CPGSV21]. matter [ABD^{+}23, ADK^{+}21, FML^{+}22, GR24, Hal20, PK21, RTK22]. measurements [GHK^{+}23]. Measuring [KCS^{+}23]. Mechanical [BB20, SBPM24, TMLS22]. mechanics [MLPM23, TH20, TH22, WA20]. mechanisms [HG21]. Medal [Hal20]. media [VPC^{+}23]. mediated [WA20]. melting [ZSC^{+}22]. Mesoscopic [BMD22]. metal [SSN^{+}23, ZSC^{+}22]. metals [HM22]. metamaterials [YZX^{+}24]. methods [LPS22, Sik21]. Micius [LCPP22]. Miniature [Sun23]. mitigation [CBB^{+}23]. Mod [GHM^{+}22, HPF^{+}24, TH22]. model [IWW24]. modeling [PD22]. Models [LPS22, LSZ^{+}20, BHMK^{+}21, CGPS22]. Modern [BK21, SZ21]. Modes [FT20, ZSC^{+}22]. molecular [PZLCEF^{+}22]. molecule [EKTvR20]. molecules [GHM^{+}18, GHM^{+}22]. momentum [JLL^{+}21]. monochalcogenide [BLFV^{+}21]. monolayers [BLFV^{+}21]. Moriya [KCS^{+}23]. motion [GBMB22]. motor [MGF24]. Multiconfigurational [LLM^{+}20]. multifractality [Jag21]. Multiple [Par23]. Multiscale [PD22]. muonic [PLH^{+}24]. mysteries [Var20].

nano [DPV^{+}20]. nano-optics [DPV^{+}20]. nanoarchitectures [PZLCEF^{+}22]. nanomechanical [BMD22]. nanoparticle [SCB^{+}23]. nanostructures [JNP^{+}23, RRM22]. Near [BMV^{+}21]. Near-field [BMV^{+}21]. network [PD22, Rei22]. networks [AEE^{+}23]. neutrino [VTR20, BBG^{+}20]. Neutrinoless [BED^{+}20, ABD^{+}23]. neutrinos [ADK^{+}21]. neutron [CSL^{+}21]. neutron-capture [CSL^{+}21]. Nobel [Gen22, Par23]. nodes [Rei22]. Noisy [BCLK^{+}22]. non [BBK21, CGPS22]. non-Fermi [CGPS22]. non-Hermitian [BBK21]. Nonequilibrium [LPS22]. nonlinear [WHK^{+}22]. nonrelativistic [SN22]. Nonthermal [dlTKC^{+}21]. Novel [BBD^{+}21]. Nuclear [HKvK20, BBG^{+}20, BDS^{+}22]. nuclei [OGS^{+}20]. NV [BSB^{+}20]. NV-diamond [BSB^{+}20].

one [BHMK^{+}21]. one-dimensional [BHMK^{+}21]. Optical [HPF^{+}22, ADKB23, RTK22, HPF^{+}24]. optics [DPV^{+}20, FT20]. optimization [BSB^{+}20]. organization [PD22, WHK^{+}22]. Origin [CSL^{+}21]. other [Hal20, MLPM23]. overview [HPB^{+}23]. oxide [JNP^{+}23].

pair [CPGSV21]. Parameter [CM22]. particle [FKK^{+}23, GBMB22, KPC^{+}20]. particles [LLM^{+}20]. Past [CXZL22]. path [HLE^{+}22]. paths [ABBW22]. pathways [dlTKC^{+}21]. perspective [LQD21]. perspectives [HKvK20, MHHT20]. phases [BG23, CGFA22, JNP^{+}23]. phenomenology [FKK^{+}23]. photoemission [SHS21]. photon [SPI^{+}23]. photon-photon [SPI^{+}23]. photonics [GVFDMM^{+}22]. photons [HLE^{+}22]. Phys [GHM^{+}22, HPF^{+}24, TH22]. Physical [BLFV^{+}21, TMNT21]. Physics [HPB^{+}23, Aer21, AAB^{+}22, BMD22, BMMS23, BDS^{+}22, GL20, Hal20, Zha23, BK21]. pinning [WKVT24]. Planckian [HM22]. plasmas [HPF^{+}22, HPF^{+}24, SBB^{+}23]. plasmon [GVFDMM^{+}22]. plasticity [BB20]. polar [JNP^{+}23]. polarized [Wu21]. polymer [LSZ^{+}20]. Positronium [BMMS23]. powder [ZSC^{+}22]. Power [Sch22, WHK^{+}22]. present [CXZL22]. principles [HPB^{+}23]. Probes [BBD^{+}21]. Probing [Aer21]. problem [ABBW22]. process [CSL^{+}21]. produced [HPF^{+}22, HPF^{+}24]. product [CPGSV21]. production [LP21]. program [HPB^{+}23]. Programmable [MCD^{+}21]. Project [BBD^{+}21]. projected [CPGSV21]. Promoting [Tho20]. properties [BLFV^{+}21, LPS22]. protected [SBPM24]. Proton [SBB^{+}23, AST21, BEG^{+}23, GV22].

QCD [BHMV21]. Quantitative [SKS^{+}23]. Quantum [AEE^{+}23, CBB^{+}23, CLM23, HLE^{+}22, MEAG^{+}20, NO20, PK21, TH20, TH22, UCNG20, ZLM^{+}23, AAB^{+}22, BHMK^{+}21, BCLK^{+}22, BGGW21, BZD24, DDM^{+}23, FT20, GHK^{+}23, HE23, KPC^{+}20, LP21, LPS22, LCPP22, MSUY22, MCD^{+}21, PZLCEF^{+}22, PR22, Rei22, SPI^{+}23, SHS21, SN22, WA20, XMZ^{+}20, ZT23, dlTKC^{+}21, MA20]. quasicrystal [Jag21]. qubits [BLP^{+}23].

radiance [SPI^{+}23]. radiation [AHM^{+}21, CLvdL23, GBMB22]. radiative [BMV^{+}21]. radio [Zha23]. radius [GV22]. random [HE23]. range [DDM^{+}23]. rapid [CSL^{+}21]. rate [CS21]. ray [GOP^{+}21]. rays [Wit20]. realistic [XMZ^{+}20]. recommended [TMNT21]. reformulations [WA20]. renormalization [MSUY22]. repeaters [AEE^{+}23]. Research [Hal20]. resistance [CXZL22]. resistivity [Var20]. resolution [MA20]. resolved [BZD24, SHS21]. resonant [GOP^{+}21]. Respectful [Tho20]. Respiratory [PPK^{+}23]. Rev [GHM^{+}22, HPF^{+}24, TH22]. Review [Kam23]. Reviews [BK21]. Role [BK21]. roles [Pic23]. Room [Pic23].

Sachdev [CGPS22]. sampling [HE23]. scalarization [DRS^{+}24]. scale [BCLK^{+}22, BB21]. scales [BBD^{+}21]. scattering [GOP^{+}21]. Science [TT22, GHK^{+}23, MLPM23]. search [Sik21]. Second [SO22]. Secure [XMZ^{+}20]. Security [PR22, BBG^{+}20]. self [HG21, WHK^{+}22]. self-limiting [HG21]. self-organization [WHK^{+}22]. Semiconductor [BLP^{+}23]. semimetals [LQD21]. Semitauonic [BSRW22]. sensitivities [WTC^{+}21]. Sensitivity [BSB^{+}20]. sensors [MA20]. shaping [HKP^{+}22]. shell [OGS^{+}20]. shift [PLH^{+}24]. sign [ABBW22]. simulations [MCD^{+}21]. single [EKTvR20, KPC^{+}20]. single-molecule [EKTvR20]. single-particle [KPC^{+}20]. singularities [Wit20]. skyrmions [RRM22]. Sliding [WKVT24]. solid [BB20]. solids [SKS^{+}23]. solution [LSZ^{+}20]. solutions [LSZ^{+}20]. Sources [VTR20]. space [LCPP22]. Specker [BCG^{+}22]. spectral [VTR20]. spectrometry [KJPW23]. spectroscopy [KPC^{+}20]. spectrum [VTR20]. spin [BLP^{+}23, MCD^{+}21, Wu21]. spin-polarized [Wu21]. Spontaneous [DRS^{+}24]. Spoof [GVFDMM^{+}22]. standard [IWW24]. stars [Aer21]. states [CPGSV21, FT20, FML^{+}22, PZLCEF^{+}22]. Statics [RRM22]. Statistical [TH20, TH22]. statistics [ADKB23]. Status [HKvK20]. steering [UCNG20]. Strong [DPV^{+}20, GBMB22, TH20, TH22]. Strong-field [DPV^{+}20]. structurally [WKVT24]. structure [KNS20, OGS^{+}20]. studies [BZD24, SHS21]. study [Jag21]. Sun [SS20]. superconductivity [Pic23, SSN^{+}23, Var20]. superfluid [NO20]. surface [GVFDMM^{+}22, PZLCEF^{+}22]. symmetries [CPGSV21]. Synchrotron [CLvdL23]. system [CS21, SO22]. systems [BMD22, BBK21, BMV^{+}21, DDM^{+}23, LPS22, MCD^{+}21, PK21, SBPM24].

techniques [CLvdL23]. technologies [AAB^{+}22]. temperature [BHMK^{+}21, Pic23, Var20]. Tensor [MSUY22]. Tests [BBD^{+}21]. their [CLvdL23]. theorem [WA20]. theorems [CPGSV21]. Theoretical [MHHT20]. theory [CGFA22, HKvK20, IWW24, JLL^{+}21, MSUY22, PLH^{+}24, Pic23, Sch22, SKS^{+}23]. thermal [CXZL22]. thermalization [BHMV21]. thermodynamics [TH20, TH22, SO22]. Thirty [MTGS20]. three [LQD21]. three-dimensional [LQD21]. Time [BZD24, LLM^{+}20, ZLM^{+}23]. time-dependent [LLM^{+}20]. Time-resolved [BZD24]. tools [BBG^{+}20, Hal20]. Topological [JNP^{+}23, LQD21]. Topologically [SBPM24]. Topology [Hal20, BBK21]. torques [FRG^{+}24]. transfer [BMV^{+}21]. transmission [PPK^{+}23]. transport [BHMK^{+}21, EKTvR20, PK21, SBPM24]. trapped [MCD^{+}21].

U.S. [HPB^{+}23]. Ultrafast [FML^{+}22, dlTKC^{+}21]. ultrathin [KCS^{+}23]. Unconventional [SSN^{+}23]. undetected [HLE^{+}22]. unification [BB21]. unified [VTR20]. universality [BSRW22]. Unusual [SS20].

values [TMNT21]. vapor [CS21]. variability [GL20]. variables [FKK^{+}23]. Vector [AST21]. Visualizing [FML^{+}22]. vortices [RTK22].

Wall [Wu21]. Water [CS21]. wave [BG23]. Waveguide [SPI^{+}23]. waves [CM22]. Weak [HHRR23, BB21]. Welcoming [TT22]. Window [CGPS22]. work [IWW24].

X [GOP^{+}21]. X-ray [GOP^{+}21].

Ye [CGPS22]. year [Gen22]. years [MTGS20].

References

Amico:2022:CAC

[AAB^{+}22] Luigi Amico, Dana Anderson, Malcolm Boshier, Jean-Philippe Brantut, Leong-Chuan Kwek, Anna Minguzzi, and Wolf von Klitzing. Colloquium: Atomtronic circuits: From manybody physics to quantum technologies. Reviews of Modern Physics, 94(4):041001–??, April 2022. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

              0756. URL http://journals.aps.org/rmp/abstract/10.
              1103/RevModPhys.94.041001.

Alexandru:2022:CPA

[ABBW22] Andrei Alexandru, Gökçe Basar, Paulo F. Bedaque, and Neill C. Warrington. Complex paths around the sign problem. Reviews of Modern Physics, 94(1):015006–??, January 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

              org/rmp/abstract/10.1103/RevModPhys.94.015006.

Agostini:2023:TDM

[ABD^{+}23] Matteo Agostini, Giovanni Benato, Jason A. Detwiler, Javier Menéndez, and Francesco Vissani. Toward the discovery of matter creation with neutrinoless ββ decay. Reviews of Modern Physics, 95(2):025002–??, February 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://www.w3.org/1998/Math/MathML.

Arguelles:2021:DMA

[ADK^{+}21] Carlos A. Argüelles, Alejandro Diaz, Ali Kheirandish, Andrés Olivares-Del-Campo, Ibrahim Safa, and Aaron C. Vincent. Dark matter annihilation to neutrinos. Reviews of Modern Physics, 93(3):035007–??, March 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.035007.

Afek:2023:CAS

[ADKB23] Gadi Afek, Nir Davidson, David A. Kessler, and Eli Barkai. Colloquium: Anomalous statistics of laser-cooled atoms in dissipative optical lattices. Reviews of Modern Physics, 95(3): 031003–??, March 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//journals.aps.org/rmp/abstract/10.1103/RevModPhys.
95.031003.

Azuma:2023:QRQ

[AEE^{+}23] Koji Azuma, Sophia E. Economou, David Elkouss, Paul Hilaire, Liang Jiang, Hoi-Kwong Lo, and Ilan Tzitrin. Quantum repeaters: From quantum networks to the quantum internet. Reviews of Modern Physics, 95(4):045006–??, April 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.95.045006.

Aerts:2021:PIP

[Aer21] C. Aerts. Probing the interior physics of stars through asteroseismology. Reviews of Modern Physics, 93(1):015001–??, January 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.93.015001.

Almheiri:2021:EHR

[AHM^{+}21] Ahmed Almheiri, Thomas Hartman, Juan Maldacena, Edgar Shaghoulian, and Amirhossein Tajdini. The entropy of Hawking radiation. Reviews of Modern Physics, 93(3):035002–??, March 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.93.035002.

Azzurri:2021:VBJ

[AST21] Paolo Azzurri, Marek Schönherr, and Alessandro Tricoli. Vector bosons and jets in proton collisions. Reviews of Modern Physics, 93(2):025007–??, February 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.025007.

Beamish:2020:MBS

[BB20] John Beamish and Sébastien Balibar. Mechanical behavior of solid helium: Elasticity, plasticity, and defects. Reviews of Modern Physics, 92(4):045002–??, April 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.92.045002.

Bodeker:2021:BWS

[BB21] Dietrich Bödeker and Wilfried Buchmüller. Baryogenesis from the weak scale to the grand unification scale. Reviews of Modern Physics, 93(3):035004–??, March 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.035004.

Baker:2021:NPP

[BBD^{+}21] Tessa Baker, Alexandre Barreira, Harry Desmond, Pedro Ferreira, Bhuvnesh Jain, Kazuya Koyama, Baojiu Li, Lucas Lombriser, Andrina Nicola, Jeremy Sakstein, and Fabian Schmidt. Novel Probes Project: Tests of gravity on astrophysical scales. Reviews of Modern Physics, 93(1):015003–??, January 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.93.015003.

Bernstein:2020:CND

[BBG^{+}20] Adam Bernstein, Nathaniel Bowden, Bethany L. Goldblum, Patrick Huber, Igor Jovanovic, and John Mattingly. Colloquium: Neutrino detectors as tools for nuclear security. Reviews of Modern Physics, 92(1):011003–??, January 2020. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.92.011003.

Bergholtz:2021:ETN

[BBK21] Emil J. Bergholtz, Jan Carl Budich, and Flore K. Kunst. Exceptional topology of non-Hermitian systems. Reviews of Modern Physics, 93(1):015005–??, January 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.93.015005.

Budroni:2022:KSC

[BCG^{+}22] Costantino Budroni, Adán Cabello, Otfried Gühne, Matthias Kleinmann, and Jan-Åke Larsson. Kochen–Specker contextuality. Reviews of Modern Physics, 94(4):045007–??, April 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.94.045007.

Bharti:2022:NIS

[BCLK^{+}22] Kishor Bharti, Alba Cervera-Lierta, Thi Ha Kyaw, Tobias Haug, Sumner Alperin-Lea, Abhinav Anand, Matthias Degroote, Hermanni Heimonen, Jakob S. Kottmann, Tim Menke, Wai-Keong Mok, Sukin Sim, Leong-Chuan Kwek, and Alán Aspuru-Guzik. Noisy intermediate-scale quantum algorithms. Reviews of Modern Physics, 94(1):015004–??, January 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.94.015004.

Blanc:2023:HHE

[BDK^{+}23] Christophe Blanc, Guillaume Durey, Randall D. Kamien, Teresa Lopez-Leon, Maxim O. Lavrentovich, and Lisa Tran. Helfrich–Hurault elastic instabilities driven by geometrical frustration. Reviews of Modern Physics, 95(1):015004–??, January 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.95.015004.

Boehnlein:2022:CML

[BDS^{+}22] Amber Boehnlein, Markus Diefenthaler, Nobuo Sato, Malachi Schram, Veronique Ziegler, Cristiano Fanelli, Morten Hjorth-Jensen, Tanja Horn, Michelle P. Kuchera, Dean Lee, Witold Nazarewicz, Peter Ostroumov, Kostas Orginos, Alan Poon, Xin-Nian Wang, Alexander Scheinker, Michael S. Smith, and Long-Gang Pang. Colloquium: Machine learning in nuclear physics. Reviews of Modern Physics, 94(3):031003–??, March 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.94.031003.

Blaum:2020:NDE

[BED^{+}20] K. Blaum, S. Eliseev, F. A. Danevich, V. I. Tretyak, Sergey Kovalenko, M. I. Krivoruchenko, Yu. N. Novikov, and J. Suhonen. Neutrinoless double-electron capture. Reviews of Modern Physics, 92(4):045007–??, April 2020. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.92.045007.

Burkert:2023:CGF

[BEG^{+}23] V. D. Burkert, L. Elouadrhiri, F. X. Girod, C. Lorcé, P. Schweitzer, and P. E. Shanahan. Colloquium: Gravitational form factors of the proton. Reviews of Modern Physics, 95(4):041002–??, April 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.041002.

Baggioli:2023:CHH

[BG23] Matteo Baggioli and Blaise Goutéraux. Colloquium: Hydrodynamics and holography of charge density wave phases. Reviews of Modern Physics, 95(1):011001–??, January 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.95.011001.

Blais:2021:CQE

[BGGW21] Alexandre Blais, Arne L. Grimsmo, S. M. Girvin, and Andreas Wallraff. Circuit quantum electrodynamics. Reviews of Modern Physics, 93(2):025005–??, February 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.93.025005.

Bertini:2021:FTT

[BHMK^{+}21] B. Bertini, F. Heidrich-Meisner, C. Karrasch, T. Prosen, R. Steinigeweg, and M. Znidaric. Finite-temperature transport in one-dimensional quantum lattice models. Reviews of Modern Physics, 93(2):025003–??, February 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.93.025003.

Berges:2021:QTI

[BHMV21] Jürgen Berges, Michal P. Heller, Aleksas Mazeliauskas, and Raju Venugopalan. QCD thermalization: Ab initio approaches and interdisciplinary connections. Reviews of Modern Physics, 93(3):035003–??, March 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.035003.

Belitz:2021:ERC

[BK21] Dietrich Belitz and Randall D. Kamien. Editorial: The role of colloquia in Reviews of Modern Physics. Reviews of Modern Physics, 93(1):010001–??, January 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.010001.

Barraza-Lopez:2021:CPP

[BLFV^{+}21] Salvador Barraza-Lopez, Benjamin M. Fregoso, John W. Villanova, Stuart S. P. Parkin, and Kai Chang. Colloquium: Physical properties of group-IV monochalcogenide monolayers. Reviews of Modern Physics, 93(1):011001–??, January 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.93.011001.

Burkard:2023:SSQ

[BLP^{+}23] Guido Burkard, Thaddeus D. Ladd, Andrew Pan, John M. Nichol, and Jason R. Petta. Semiconductor spin qubits. Reviews of Modern Physics, 95(2):025003–??, February 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.95.025003.

Bachtold:2022:MPN

[BMD22] Adrian Bachtold, Joel Moser, and M. I. Dykman. Mesoscopic physics of nanomechanical systems. Reviews of Modern Physics, 94(4):045005–??, April 2022. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.94.045005.

Bass:2023:CPP

[BMMS23] Steven D. Bass, Sebastiano Mariazzi, Pawel Moskal, and Ewa Stepień. Colloquium: Positronium physics and biomedical applications. Reviews of Modern Physics, 95(2):021002–??, February 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.95.021002.

Biehs:2021:NFR

[BMV^{+}21] S.-A. Biehs, R. Messina, P. S. Venkataram, A. W. Rodriguez, J. C. Cuevas, and P. Ben-Abdallah. Near-field radiative heat transfer in many-body systems. Reviews of Modern Physics, 93(2):025009–??, February 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.025009.

Barry:2020:SON

[BSB^{+}20] John F. Barry, Jennifer M. Schloss, Erik Bauch, Matthew J. Turner, Connor A. Hart, Linh M. Pham, and Ronald L. Walsworth. Sensitivity optimization for NV-diamond magnetometry. Reviews of Modern Physics, 92(1):015004–??, January 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.92.015004.

Bernlochner:2022:SHD

[BSRW22] Florian U. Bernlochner, Manuel Franco Sevilla, Dean J. Robinson, and Guy Wormser. Semitauonic b-hadron decays: A lepton flavor universality laboratory. Reviews of Modern Physics, 94(1):015003–??, January 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://www.w3.org/1998/Math/MathML.

Boschini:2024:TRA

[BZD24] Fabio Boschini, Marta Zonno, and Andrea Damascelli. Timeresolved ARPES studies of quantum materials. Reviews of Modern Physics, 96(1):015003–??, January 2024. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.96.015003.

Cai:2023:QEM

[CBB^{+}23] Zhenyu Cai, Ryan Babbush, Simon C. Benjamin, Suguru Endo, William J. Huggins, Ying Li, Jarrod R. McClean, and Thomas E. O’Brien. Quantum error mitigation. Reviews of Modern Physics, 95(4):045005–??, April 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.045005.

Cisowski:2022:CGP

[CGFA22] C. Cisowski, J. B. Götte, and S. Franke-Arnold. Colloquium: Geometric phases of light: Insights from fiber bundle theory. Reviews of Modern Physics, 94(3):031001–??, March 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.94.031001.

Chowdhury:2022:SYK

[CGPS22] Debanjan Chowdhury, Antoine Georges, Olivier Parcollet, and Subir Sachdev. Sachdev–Ye–Kitaev models and beyond: Window into non-Fermi liquids. Reviews of Modern Physics, 94(3):035004–??, March 2022. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.94.035004.

Chang:2023:CQA

[CLM23] Cui-Zu Chang, Chao-Xing Liu, and Allan H. MacDonald. Colloquium: Quantum anomalous Hall effect. Reviews of Modern Physics, 95(1):011002–??, January 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.011002.

Caciuffo:2023:SRT

[CLvdL23] R. Caciuffo, G. H. Lander, and G. van der Laan. Synchrotron radiation techniques and their application to actinide materials. Reviews of Modern Physics, 95(1):015001–??, January 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.95.015001.

Christensen:2022:PEG

[CM22] Nelson Christensen and Renate Meyer. Parameter estimation with gravitational waves. Reviews of Modern Physics, 94(2):025001–??, February 2022. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.94.025001.

Cirac:2021:MPS

[CPGSV21] J. Ignacio Cirac, David Pérez-García, Norbert Schuch, and Frank Verstraete. Matrix product states and projected entangled pair states: Concepts, symmetries, theorems. Reviews of Modern Physics, 93(4):045003–??, April 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.93.045003.

Colman:2021:WVL

[CS21] Robert Colman and Brian J. Soden. Water vapor and lapse rate feedbacks in the climate system. Reviews of Modern Physics, 93(4):045002–??, April 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.045002.

Cowan:2021:OHE

[CSL^{+}21] John J. Cowan, Christopher Sneden, James E. Lawler, Ani Aprahamian, Michael Wiescher, Karlheinz Langanke, Gabriel Martínez-Pinedo, and Friedrich-Karl Thielemann. Origin of the heaviest elements: The rapid neutron-capture process. Reviews of Modern Physics, 93(1):015002–??, January 2021. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.93.015002.

Chen:2022:ITR

[CXZL22] Jie Chen, Xiangfan Xu, Jun Zhou, and Baowen Li. Interfacial thermal resistance: Past, present, and future. Reviews of Modern Physics, 94(2):025002–??, February 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.94.025002.

Defenu:2023:LRI

[DDM^{+}23] Nicolò Defenu, Tobias Donner, Tommaso Macrì, Guido Pagano, Stefano Ruffo, and Andrea Trombettoni. Long-range interacting quantum systems. Reviews of Modern Physics, 95 (3):035002–??, March 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//journals.aps.org/rmp/abstract/10.1103/RevModPhys.
95.035002.

delaTorre:2021:CNP

[dlTKC^{+}21] Alberto de la Torre, Dante M. Kennes, Martin Claassen, Simon Gerber, James W. McIver, and Michael A. Sentef. Colloquium: Nonthermal pathways to ultrafast control in quantum materials. Reviews of Modern Physics, 93(4):041002–??, April 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.93.041002.

Dombi:2020:SFN

[DPV^{+}20] Péter Dombi, Zsuzsanna Pápa, Jan Vogelsang, Sergey V. Yalunin, Murat Sivis, Georg Herink, Sascha Schäfer, Petra Groß, Claus Ropers, and Christoph Lienau. Strong-field nanooptics. Reviews of Modern Physics, 92(2):025003–??, February 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.92.025003.

Doneva:2024:SS

[DRS^{+}24] Daniela D. Doneva, Fethi M. Ramazanoglu, Hector O. Silva, Thomas P. Sotiriou, and Stoytcho S. Yazadjiev. Spontaneous scalarization. Reviews of Modern Physics, 96(1):015004–??, January 2024. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.96.015004.

Emparan:2020:LLE

[EH20] Roberto Emparan and Christopher P. Herzog. Large D limit of Einstein’s equations. Reviews of Modern Physics, 92(4): 045005–??, April 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://www.

w3.org/1998/Math/MathML.

Evers:2020:ACS

[EKTvR20] Ferdinand Evers, Richard Korytár, Sumit Tewari, and Jan M. van Ruitenbeek. Advances and challenges in single-molecule electron transport. Reviews of Modern Physics, 92(3):035001– ??, March 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.92.035001.

Franceschini:2023:KVF

[FKK^{+}23] Roberto Franceschini, Doojin Kim, Kyoungchul Kong, Konstantin T. Matchev, Myeonghun Park, and Prasanth Shyamsundar. Kinematic variables and feature engineering for particle phenomenology. Reviews of Modern Physics, 95(4):045004– ??, April 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.95.045004.

Filippetto:2022:UED

[FML^{+}22] D. Filippetto, P. Musumeci, R. K. Li, B. J. Siwick, M. R. Otto, M. Centurion, and J. P. F. Nunes. Ultrafast electron diffraction: Visualizing dynamic states of matter. Reviews of Modern Physics, 94(4):045004–??, April 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.94.045004.

Fert:2024:ECM

[FRG^{+}24] Albert Fert, Ramamoorthy Ramesh, Vincent Garcia, Fèlix Casanova, and Manuel Bibes. Electrical control of magnetism by electric field and current-induced torques. Reviews of Modern Physics, 96(1):015005–??, January 2024. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.96.015005.

Fabre:2020:MSQ

[FT20] C. Fabre and N. Treps. Modes and states in quantum optics. Reviews of Modern Physics, 92(3):035005–??, March 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.92.035005.

Gonoskov:2022:CPM

[GBMB22] A. Gonoskov, T. G. Blackburn, M. Marklund, and S. S. Bulanov. Charged particle motion and radiation in strong electromagnetic fields. Reviews of Modern Physics, 94(4):045001– ??, April 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.94.045001.

Genzel:2022:NLF

[Gen22] Reinhard Genzel. Nobel Lecture: A forty-year journey. Reviews of Modern Physics, 94(2):020501–??, February 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.94.020501.

Guhne:2023:CIM

[GHK^{+}23] Otfried Gühne, Erkka Haapasalo, Tristan Kraft, Juha-Pekka Pellonpää, and Roope Uola. Colloquium: Incompatible measurements in quantum information science. Reviews of Modern Physics, 95(1):011003–??, January 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.011003.

Guo:2018:HM

[GHM^{+}18] Feng-Kun Guo, Christoph Hanhart, Ulf-G. Meißner, Qian Wang, Qiang Zhao, and Bing-Song Zou. Hadronic molecules. Reviews of Modern Physics, 90(1):015004–??, January 2018. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.90.015004. See erra-

tum [GHM^{+}22].

Guo:2022:EHM

[GHM^{+}22] Feng-Kun Guo, Christoph Hanhart, Ulf-G. Meißner, Qian Wang, Qiang Zhao, and Bing-Song Zou. Erratum: Hadronic molecules [Rev. Mod. Phys. 90, 015004 (2018)]. Reviews of Modern Physics, 94(2):029901–??, February 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.94.029901. See [GHM^{+}18].

Ghil:2020:PCV

[GL20] Michael Ghil and Valerio Lucarini. The physics of climate variability and climate change. Reviews of Modern Physics, 92(3):035002–??, March 2020. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.92.035002.

Gelmukhanov:2021:DRX

[GOP^{+}21] Faris Gel’mukhanov, Michael Odelius, Sergey P. Polyutov, Alexander Föhlisch, and Victor Kimberg. Dynamics of resonant X-ray and Auger scattering. Reviews of Modern Physics, 93(3):035001–??, March 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.035001.

Gromov:2024:CFM

[GR24] Andrey Gromov and Leo Radzihovsky. Colloquium: Fracton matter. Reviews of Modern Physics, 96(1):011001–??, January 2024. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.96.011001.

Gao:2022:PCR

[GV22] H. Gao and M. Vanderhaeghen. The proton charge radius. Reviews of Modern Physics, 94(1):015002–??, January 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.94.015002.

Garcia-Vidal:2022:SSP

[GVFDMM^{+}22] Francisco J. Garcia-Vidal, Antonio I. Fernández-Domínguez, Luis Martin-Moreno, Hao Chi Zhang, Wenxuan Tang, Ruwen Peng, and Tie Jun Cui. Spoof surface plasmon photonics. Reviews of Modern Physics, 94(2):025004–??, February 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.94.025004.

Halperin:2020:AME

[Hal20] Bertrand I. Halperin. APS Medal for Exceptional Achievement in Research: Topology and other tools in condensed matter physics. Reviews of Modern Physics, 92(4):045001– ??, April 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.92.045001.

Hangleiter:2023:CAQ

[HE23] Dominik Hangleiter and Jens Eisert. Computational advantage of quantum random sampling. Reviews of Modern Physics, 95(3):035001–??, March 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.035001.

Hagan:2021:EMS

[HG21] Michael F. Hagan and Gregory M. Grason. Equilibrium mechanisms of self-limiting assembly. Reviews of Modern Physics, 93(2):025008–??, February 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.025008.

Harlow:2023:WGC

[HHRR23] Daniel Harlow, Ben Heidenreich, Matthew Reece, and Tom Rudelius. Weak gravity conjecture. Reviews of Modern Physics, 95(3):035003–??, March 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.035003.

Ha:2022:BSE

[HKP^{+}22] G. Ha, K.-J. Kim, J. G. Power, Y. Sun, and P. Piot. Bunch shaping in electron linear accelerators. Reviews of Modern Physics, 94(2):025006–??, February 2022. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.94.025006.

Hammer:2020:NEF

[HKvK20] H.-W. Hammer, Sebastian König, and U. van Kolck. Nuclear effective field theory: Status and perspectives. Reviews of Modern Physics, 92(2):025004–??, February 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.92.025004.

Hochrainer:2022:QIP

[HLE^{+}22] Armin Hochrainer, Mayukh Lahiri, Manuel Erhard, Mario Krenn, and Anton Zeilinger. Quantum indistinguishability by path identity and with undetected photons. Reviews of Modern Physics, 94(2):025007–??, February 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.94.025007.

Hartnoll:2022:CPD

[HM22] Sean A. Hartnoll and Andrew P. Mackenzie. Colloquium: Planckian dissipation in metals. Reviews of Modern Physics, 94(4):041002–??, April 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//journals.aps.org/rmp/abstract/10.1103/RevModPhys.
94.041002.

Hurricane:2023:PPI

[HPB^{+}23] O. A. Hurricane, P. K. Patel, R. Betti, D. H. Froula, S. P. Regan, S. A. Slutz, M. R. Gomez, and M. A. Sweeney. Physics principles of inertial confinement fusion and U.S. program overview. Reviews of Modern Physics, 95(2):025005–??, February 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.95.025005.

Harilal:2022:ODL

[HPF^{+}22] S. S. Harilal, M. C. Phillips, D. H. Froula, K. K. Anoop, R. C. Issac, and F. N. Beg. Optical diagnostics of laser-produced plasmas. Reviews of Modern Physics, 94(3):035002–??, March 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.94.035002. See erra-

tum [HPF^{+}24].

Harilal:2024:EOD

[HPF^{+}24] S. S. Harilal, M. C. Phillips, D. H. Froula, K. K. Anoop, R. C. Issac, and F. N. Beg. Erratum: Optical diagnostics of laser-produced plasmas [Rev. Mod. Phys. 94, 035002 (2022)]. Reviews of Modern Physics, 96(1):019901–??, January 2024. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.96.019901. See [HPF^{+}22].

Isidori:2024:SME

[IWW24] Gino Isidori, Felix Wilsch, and Daniel Wyler. The standard model effective field theory at work. Reviews of Modern Physics, 96(1):015006–??, January 2024. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.96.015006.

Jagannathan:2021:FQC

[Jag21] Anuradha Jagannathan. The Fibonacci quasicrystal: Case study of hidden dimensions and multifractality. Reviews of Modern Physics, 93(4):045001–??, April 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.045001.

Ji:2021:LME

[JLL^{+}21] Xiangdong Ji, Yizhuang Liu, Yu-Sheng Liu, Jian-Hui Zhang, and Yong Zhao. Large-momentum effective theory. Reviews of Modern Physics, 93(3):035005–??, March 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.93.035005.

Junquera:2023:TPP

[JNP^{+}23] Javier Junquera, Yousra Nahas, Sergei Prokhorenko, Laurent Bellaiche, Jorge Íñiguez, Darrell G. Schlom, Long-Qing Chen, Sayeef Salahuddin, David A. Muller, Lane W. Martin, and R. Ramesh. Topological phases in polar oxide nanostructures. Reviews of Modern Physics, 95(2):025001–??, February 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.95.025001.

Kamien:2023:ER

[Kam23] Randall D. Kamien. Editorial: To review is to be. Reviews of Modern Physics, 95(3):030001–??, March 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.95.030001.

Kuepferling:2023:MID

[KCS^{+}23] M. Kuepferling, A. Casiraghi, G. Soares, G. Durin, F. Garcia-Sanchez, L. Chen, C. H. Back, C. H. Marrows, S. Tacchi, and G. Carlotti. Measuring interfacial Dzyaloshinskii–Moriya interaction in ultrathin magnetic films. Reviews of Modern Physics, 95(1):015003–??, January 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.015003.

Kutschera:2023:ACA

[KJPW23] Walter Kutschera, A. J. Timothy Jull, Michael Paul, and Anton Wallner. Atom counting with accelerator mass spectrometry. Reviews of Modern Physics, 95(3):035006–??, March 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.95.035006.

Kovarik:2020:HSH

[KNS20] Karol Kovarík, Pavel M. Nadolsky, and Davison E. Soper. Hadronic structure in high-energy collisions. Reviews of Modern Physics, 92(4):045003–??, April 2020. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.92.045003.

Kirchner:2020:CHE

[KPC^{+}20] Stefan Kirchner, Silke Paschen, Qiuyun Chen, Steffen Wirth, Donglai Feng, Joe D. Thompson, and Qimiao Si. Colloquium: Heavy-electron quantum criticality and single-particle spectroscopy. Reviews of Modern Physics, 92(1):011002–??, January 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.92.011002.

Lu:2022:MQE

[LCPP22] Chao-Yang Lu, Yuan Cao, Cheng-Zhi Peng, and Jian-Wei Pan. Micius quantum experiments in space. Reviews of Modern Physics, 94(3):035001–??, March 2022. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.94.035001.

Lode:2020:CMT

[LLM^{+}20] Axel U. J. Lode, Camille Lévêque, Lars Bojer Madsen, Alexej I. Streltsov, and Ofir E. Alon. Colloquium: Multiconfigurational time-dependent Hartree approaches for indistinguishable particles. Reviews of Modern Physics, 92(1):011001– ??, January 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.92.011001.

Landi:2021:IEP

[LP21] Gabriel T. Landi and Mauro Paternostro. Irreversible entropy production: From classical to quantum. Reviews of Modern Physics, 93(3):035008–??, March 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.035008.

Landi:2022:NBD

[LPS22] Gabriel T. Landi, Dario Poletti, and Gernot Schaller. Nonequilibrium boundary-driven quantum systems: Models, methods, and properties. Reviews of Modern Physics, 94(4):045006– ??, April 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.94.045006.

Lv:2021:EPT

[LQD21] B. Q. Lv, T. Qian, and H. Ding. Experimental perspective on three-dimensional topological semimetals. Reviews of Modern Physics, 93(2):025002–??, February 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.93.025002.

Lauricella:2020:MPS

[LSZ^{+}20] Marco Lauricella, Sauro Succi, Eyal Zussman, Dario Pisignano, and Alexander L. Yarin. Models of polymer solutions in electrified jets and solution blowing. Reviews of Modern Physics, 92(3):035004–??, March 2020. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.92.035004.

Mitchell:2020:CQL

[MA20] Morgan W. Mitchell and Silvana Palacios Alvarez. Colloquium: Quantum limits to the energy resolution of magnetic field sensors. Reviews of Modern Physics, 92(2):021001–??, February 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.92.021001.

Monroe:2021:PQS

[MCD^{+}21] C. Monroe, W. C. Campbell, L.-M. Duan, Z.-X. Gong, A. V. Gorshkov, P. W. Hess, R. Islam, K. Kim, N. M. Linke, G. Pagano, P. Richerme, C. Senko, and N. Y. Yao. Programmable quantum simulations of spin systems with trapped ions. Reviews of Modern Physics, 93(2):025001–??, February 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.93.025001.

McArdle:2020:QCC

[MEAG^{+}20] Sam McArdle, Suguru Endo, Alán Aspuru-Guzik, Simon C. Benjamin, and Xiao Yuan. Quantum computational chemistry. Reviews of Modern Physics, 92(1):015003–??, January 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.92.015003.

Marmol:2024:CMB

[MGF24] M. Marmol, E. Gachon, and D. Faivre. Colloquium: Magnetotactic bacteria: From flagellar motor to collective effects. Reviews of Modern Physics, 96(2):021001–??, February 2024. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.96.021001.

Mugnai:2020:TPB

[MHHT20] Mauro L. Mugnai, Changbong Hyeon, Michael Hinczewski, and D. Thirumalai. Theoretical perspectives on biological machines. Reviews of Modern Physics, 92(2):025001–??, February 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.92.025001.

Mathijssen:2023:CFM

[MLPM23] Arnold J. T. M. Mathijssen, Maciej Lisicki, Vivek N. Prakash, and Endre J. L. Mossige. Culinary fluid mechanics and other currents in food science. Reviews of Modern Physics, 95(2):025004–??, February 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.025004.

Meurice:2022:TLF

[MSUY22] Yannick Meurice, Ryo Sakai, and Judah Unmuth-Yockey. Tensor lattice field theory for renormalization and quantum computing. Reviews of Modern Physics, 94(2):025005–??, February 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.94.025005.

Miller:2020:TYH

[MTGS20] Steve Miller, Jonathan Tennyson, Thomas R. Geballe, and Tom Stallard. Thirty years of H_{3}^{+} astronomy. Reviews of Modern Physics, 92(3):035003–??, March 2020. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://www.w3.org/1998/Math/MathML.

Nomura:2020:CQC

[NO20] R. Nomura and Y. Okuda. Colloquium: Quantum crystallizations of ^{4}He in superfluid far from equilibrium. Reviews of Modern Physics, 92(4):041003–??, April 2020. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://www.w3.org/1998/Math/MathML.

Otsuka:2020:ESS

[OGS^{+}20] Takaharu Otsuka, Alexandra Gade, Olivier Sorlin, Toshio Suzuki, and Yutaka Utsuno. Evolution of shell structure in exotic nuclei. Reviews of Modern Physics, 92(1):015002–??, January 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.92.015002.

Parisi:2023:NLM

[Par23] Giorgio Parisi. Nobel lecture: Multiple equilibria. Reviews of Modern Physics, 95(3):030501–??, March 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.95.030501.

Presigny:2022:CMM

[PD22] Charley Presigny and Fabrizio De Vico Fallani. Colloquium: Multiscale modeling of brain network organization. Reviews of Modern Physics, 94(3):031002–??, March 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.94.031002.

Pickett:2023:CRT

[Pic23] Warren E. Pickett. Colloquium: Room temperature superconductivity: The roles of theory and materials design. Reviews of Modern Physics, 95(2):021001–??, February 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.95.021001.

Pekola:2021:CQH

[PK21] Jukka P. Pekola and Bayan Karimi. Colloquium: Quantum heat transport in condensed matter systems. Reviews of Modern Physics, 93(4):041001–??, April 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.93.041001.

Pachucki:2024:CTL

[PLH^{+}24] K. Pachucki, V. Lensky, F. Hagelstein, S. S. Li Muli, S. Bacca, and R. Pohl. Comprehensive theory of the Lamb shift in light muonic atoms. Reviews of Modern Physics, 96(1):015001–??, January 2024. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.96.015001.

Pohlker:2023:RAD

[PPK^{+}23] Mira L. Pöhlker, Christopher Pöhlker, Ovid O. Krüger, Jan-David Förster, Thomas Berkemeier, Wolfgang Elbert, Janine Fröhlich-Nowoisky, Ulrich Pöschl, Gholamhossein Bagheri, Eberhard Bodenschatz, J. Alex Huffman, Simone Scheithauer, and Eugene Mikhailov. Respiratory aerosols and droplets in the transmission of infectious diseases. Reviews of Modern Physics, 95(4):045001–??, April 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.045001.

Portmann:2022:SQC

[PR22] Christopher Portmann and Renato Renner. Security in quantum cryptography. Reviews of Modern Physics, 94(2): 025008–??, February 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//journals.aps.org/rmp/abstract/10.1103/RevModPhys.
94.025008.

Piquero-Zulaica:2022:EQS

[PZLCEF^{+}22] Ignacio Piquero-Zulaica, Jorge Lobo-Checa, Zakaria M. Abd El-Fattah, J. Enrique Ortega, Florian Klappenberger, Willi Auwärter, and Johannes V. Barth. Engineering quantum states and electronic landscapes through surface molecular nanoarchitectures. Reviews of Modern Physics, 94(4):045008– ??, April 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.94.045008.

Reiserer:2022:CCE

[Rei22] Andreas Reiserer. Colloquium: Cavity-enhanced quantum network nodes. Reviews of Modern Physics, 94(4):041003–??, April 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.94.041003.

Reichhardt:2022:SDS

[RRM22] C. Reichhardt, C. J. O. Reichhardt, and M. V. Milosević. Statics and dynamics of skyrmions interacting with disorder and nanostructures. Reviews of Modern Physics, 94(3):035005–??, March 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.94.035005.

Rosen:2022:IBO

[RTK22] Guillermo F. Quinteiro Rosen, Pablo I. Tamborenea, and Tilmann Kuhn. Interplay between optical vortices and condensed matter. Reviews of Modern Physics, 94(3):035003–??, March 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.94.035003.

Schaeffer:2023:PIH

[SBB^{+}23] Derek B. Schaeffer, Archie F. A. Bott, Marco Borghesi, Kirk A. Flippo, William Fox, Julien Fuchs, Chikang Li, Fredrick H. Séguin, Hye-Sook Park, Petros Tzeferacos, and Louise Willingale. Proton imaging of high-energy-density laboratory plasmas. Reviews of Modern Physics, 95(4):045007–??, April 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.95.045007.

Shah:2024:CTP

[SBPM24] Tirth Shah, Christian Brendel, Vittorio Peano, and Florian Marquardt. Colloquium: Topologically protected transport in engineered mechanical systems. Reviews of Modern Physics, 96(2):021002–??, February 2024. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.96.021002.

Stolzenburg:2023:ANG

[SCB^{+}23] Dominik Stolzenburg, Runlong Cai, Sara M. Blichner, Jenni Kontkanen, Putian Zhou, Risto Makkonen, Veli-Matti Kerminen, Markku Kulmala, Ilona Riipinen, and Juha Kangasluoma. Atmospheric nanoparticle growth. Reviews of Modern Physics, 95(4):045002–??, April 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.045002.

Schmidt:2022:PFT

[Sch22] Matthias Schmidt. Power functional theory for many-body dynamics. Reviews of Modern Physics, 94(1):015007–??, January 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.94.015007.

Sobota:2021:ARP

[SHS21] Jonathan A. Sobota, Yu He, and Zhi-Xun Shen. Angle-resolved photoemission studies of quantum materials. Reviews of Modern Physics, 93(2):025006–??, February 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.025006.

Sikivie:2021:IAS

[Sik21] Pierre Sikivie. Invisible axion search methods. Reviews of Modern Physics, 93(1):015004–??, January 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.93.015004.

Szilva:2023:QTM

[SKS^{+}23] Attila Szilva, Yaroslav Kvashnin, Evgeny A. Stepanov, Lars Nordström, Olle Eriksson, Alexander I. Lichtenstein, and Mikhail I. Katsnelson. Quantitative theory of magnetic interactions in solids. Reviews of Modern Physics, 95(3):035004– ??, March 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.95.035004.

Stokes:2022:IGF

[SN22] Adam Stokes and Ahsan Nazir. Implications of gauge freedom for nonrelativistic quantum electrodynamics. Reviews of Modern Physics, 94(4):045003–??, April 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.94.045003.

Singh:2022:CSS

[SO22] Martin S. Singh and Morgan E O’Neill. The climate system and the Second Law of Thermodynamics. Reviews of Modern Physics, 94(1):015001–??, January 2022. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.94.015001.

Sheremet:2023:WQE

[SPI^{+}23] Alexandra S. Sheremet, Mihail I. Petrov, Ivan V. Iorsh, Alexander V. Poshakinskiy, and Alexander N. Poddubny. Waveguide quantum electrodynamics: Collective radiance and photon-photon correlations. Reviews of Modern Physics, 95 (1):015002–??, January 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//journals.aps.org/rmp/abstract/10.1103/RevModPhys.
95.015002.

Schumacher:2020:CUD

[SS20] Jörg Schumacher and Katepalli R. Sreenivasan. Colloquium: Unusual dynamics of convection in the Sun. Reviews of Modern Physics, 92(4):041001–??, April 2020. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.92.041001.

Smidman:2023:CUF

[SSN^{+}23] Michael Smidman, Oliver Stockert, Emilian M. Nica, Yang Liu, Huiqiu Yuan, Qimiao Si, and Frank Steglich. Colloquium: Unconventional fully gapped superconductivity in the heavyfermion metal CeCu_{2}Si_{2}. Reviews of Modern Physics, 95(3): 031002–??, March 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://www.

w3.org/1998/Math/MathML.

Sun:2023:CMI

[Sun23] Mao Sun. Colloquium: Miniature insect flight. Reviews of Modern Physics, 95(4):041001–??, April 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.041001.

Shiltsev:2021:MFC

[SZ21] V. Shiltsev and F. Zimmermann. Modern and future colliders. Reviews of Modern Physics, 93(1):015006–??, January 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.93.015006.

Talkner:2020:CSM

[TH20] Peter Talkner and Peter Hänggi. Colloquium: Statistical mechanics and thermodynamics at strong coupling: Quantum and classical. Reviews of Modern Physics, 92(4):041002–??, April 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.92.041002. See

erratum [TH22].

Talkner:2022:ECS

[TH22] Peter Talkner and Peter Hänggi. Erratum: Colloquium: Statistical mechanics and thermodynamics at strong coupling: Quantum and classical [Rev. Mod. Phys. 92, 041002 (2020)]. Reviews of Modern Physics, 94(3):039901–??, March 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.94.039901. See [TH20].

Thoennessen:2020:EPI

[Tho20] Michael Thoennessen. Editorial: Promoting inclusive and respectful communications. Reviews of Modern Physics, 92 (4):040001–??, April 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//journals.aps.org/rmp/abstract/10.1103/RevModPhys.
92.040001.

Torbati:2022:CMD

[TMLS22] Mehdi Torbati, Kosar Mozaffari, Liping Liu, and Pradeep Sharma. Coupling of mechanical deformation and electromagnetic fields in biological cells. Reviews of Modern Physics, 94(2):025003–??, February 2022. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.94.025003.

Tiesinga:2021:CRV

[TMNT21] Eite Tiesinga, Peter J. Mohr, David B. Newell, and Barry N. Taylor. CODATA recommended values of the fundamental physical constants: 2018. Reviews of Modern Physics, 93(2): 025010–??, February 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//journals.aps.org/rmp/abstract/10.1103/RevModPhys.
93.025010.

Thomas:2022:EWH

[TT22] Jessica Thomas and Michael Thoennessen. Editorial: A welcoming home for applied science. Reviews of Modern Physics, 94(4):040001–??, April 2022. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.94.040001.

Uola:2020:QS

[UCNG20] Roope Uola, Ana C. S. Costa, H. Chau Nguyen, and Otfried Gühne. Quantum steering. Reviews of Modern Physics, 92(1):015001–??, January 2020. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.92.015001.

Varma:2020:CLT

[Var20] Chandra M. Varma. Colloquium: Linear in temperature resistivity and associated mysteries including high temperature superconductivity. Reviews of Modern Physics, 92(3):031001– ??, March 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.92.031001.

Vynck:2023:LCD

[VPC^{+}23] Kevin Vynck, Romain Pierrat, Rémi Carminati, Luis S. Froufe-Pérez, Frank Scheffold, Riccardo Sapienza, Silvia Vignolini, and Juan José Sáenz. Light in correlated disordered media. Reviews of Modern Physics, 95(4):045003–??, April 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.95.045003.

Vitagliano:2020:GUN

[VTR20] Edoardo Vitagliano, Irene Tamborra, and Georg Raffelt. Grand unified neutrino spectrum at Earth: Sources and spectral components. Reviews of Modern Physics, 92(4):045006– ??, April 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.92.045006.

Wharton:2020:CBT

[WA20] K. B. Wharton and N. Argaman. Colloquium: Bell’s theorem and locally mediated reformulations of quantum mechanics. Reviews of Modern Physics, 92(2):021002–??, February 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.

org/rmp/abstract/10.1103/RevModPhys.92.021002.

Witthaut:2022:CND

[WHK^{+}22] Dirk Witthaut, Frank Hellmann, Jürgen Kurths, Stefan Kettemann, Hildegard Meyer-Ortmanns, and Marc Timme. Collective nonlinear dynamics and self-organization in decentralized power grids. Reviews of Modern Physics, 94(1):015005–??, January 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.94.015005.

Witten:2020:LRS

[Wit20] Edward Witten. Light rays, singularities, and all that. Reviews of Modern Physics, 92(4):045004–??, April 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.92.045004.

Wang:2024:CSP

[WKVT24] Jin Wang, Ali Khosravi, Andrea Vanossi, and Erio Tosatti. Colloquium: Sliding and pinning in structurally lubric 2D material interfaces. Reviews of Modern Physics, 96(1):011002–??, January 2024. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.96.011002.

Williamson:2021:ECC

[WTC^{+}21] Mark S. Williamson, Chad W. Thackeray, Peter M. Cox, Alex Hall, Chris Huntingford, and Femke J. M. M. Nijsse. Emergent constraints on climate sensitivities. Reviews of Modern Physics, 93(2):025004–??, February 2021. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.93.025004.

Wu:2021:WIS

[Wu21] Zhen Wu. Wall interactions of spin-polarized atoms. Reviews of Modern Physics, 93(3):035006–??, March 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.93.035006.

Xu:2020:SQK

[XMZ^{+}20] Feihu Xu, Xiongfeng Ma, Qiang Zhang, Hoi-Kwong Lo, and Jian-Wei Pan. Secure quantum key distribution with realistic devices. Reviews of Modern Physics, 92(2):025002–??, February 2020. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.

aps.org/rmp/abstract/10.1103/RevModPhys.92.025002.

Yang:2024:CME

[YZX^{+}24] Fubao Yang, Zeren Zhang, Liujun Xu, Zhoufei Liu, Peng Jin, Pengfei Zhuang, Min Lei, Jinrong Liu, Jian-Hua Jiang, Xiaoping Ouyang, Fabio Marchesoni, and Jiping Huang. Controlling mass and energy diffusion with metamaterials. Reviews of Modern Physics, 96(1):015002–??, January 2024. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.96.015002.

Zhang:2023:PFR

[Zha23] Bing Zhang. The physics of fast radio bursts. Reviews of Modern Physics, 95(3):035005–??, March 2023. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.
1103/RevModPhys.95.035005.

Zaletel:2023:CQC

[ZLM^{+}23] Michael P. Zaletel, Mikhail Lukin, Christopher Monroe, Chetan Nayak, Frank Wilczek, and Norman Y. Yao. Colloquium: Quantum and classical discrete time crystals. Reviews of Modern Physics, 95(3):031001–??, March 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.95.031001.

Zhao:2022:LMM

[ZSC^{+}22] Cang Zhao, Bo Shi, Shuailei Chen, Dong Du, Tao Sun, Brian J. Simonds, Kamel Fezzaa, and Anthony D. Rollett. Laser melting modes in metal powder bed fusion additive manufacturing. Reviews of Modern Physics, 94(4):045002–??, April 2022. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.94.045002.

Zwolak:2023:CAA

[ZT23] Justyna P. Zwolak and Jacob M. Taylor. Colloquium: Advances in automation of quantum dot devices control. Reviews of Modern Physics, 95(1):011006–??, January 2023. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.95.011006.